OpenStudy (anonymous):

find the unique subgroup of U(11) of order 5 and list its elements?

OpenStudy (anonymous):

@mahmit2012

OpenStudy (anonymous):

i have found that 3, 4, 5 , and 9 has order 5 but what does it mean by unique subgroup and list the elements

OpenStudy (zarkon):

the other elements are not of order 5 correct?

OpenStudy (anonymous):

nope 1 has order 1 , (2,6,7,8) has order 10 and those are the generators, (3,4,5,9) have order 5 and (10) has order 2

OpenStudy (zarkon):

and 3,4,5,9 by themselves all generate the same group correct?

OpenStudy (anonymous):

uhh what do you mean by that

OpenStudy (zarkon):

the element 3 generates {1,3,4,5,9}

OpenStudy (zarkon):

the element 4 generates {1,3,4,5,9}

OpenStudy (zarkon):

...

OpenStudy (anonymous):

i am confused by that how did you get those numbers :S

OpenStudy (zarkon):

get which numbers?

OpenStudy (anonymous):

where you said the element 3 generates {1,3,4,5,9}

OpenStudy (anonymous):

i dont know what i am trying to find when it says find the unique subgroup of U(11) or order 5 and list its elements

OpenStudy (zarkon):

3 3*3=9 3*9=5 mod 11 3*5=4 mod 11 3*4=1 mod 11

OpenStudy (zarkon):

{3,9,5,4,1}

OpenStudy (zarkon):

it is the unique subgroup of order 5

OpenStudy (anonymous):

im sorry but i am still kind of confused a little ok so you know how i got (3,4,5,9) and it has order 5 right now are we picking any number in that set?

OpenStudy (anonymous):

or are we going to do it for 4 5 and 9 as well

OpenStudy (zarkon):

any number from that set generates the same group and all the other numbers beside 1 generates all of U(11)

OpenStudy (anonymous):

alright so the unique subgroup of U(11) of order 5 are (3,4,5,9) ?

OpenStudy (zarkon):

the unique subgroup of order 5 is {1,3,4,5,9}

OpenStudy (zarkon):

1 is in the group

OpenStudy (anonymous):

why are we including 1 though doesnt 1 have order 1

OpenStudy (zarkon):

you are looking at the order of the group not the order of each of the elements.

OpenStudy (anonymous):

oooo yeah yeah sorry ok

OpenStudy (zarkon):

if you don't include 1 then your 'group' is not closed under the operation

OpenStudy (anonymous):

your right my mistake